It is January 1, 2019. Find the present value of a $100 perpetuity when the first payment is at the end of year 2019. The discount rate is 6%.

It is January 1, 2019. Find the present value of a $100 perpetuity when the first payment is at the end of year 2020. The discount rate is 6%.

It is January 1, 2019. Find the present value of a $100 growing perpetuity when the first payment is at the end of year 2019. The discount rate is 6%. The growth rate is 2%
Answer #1

Present Value of $100 Perpetuity at end of year = $100/0.06 = $1666.67
Present Value on Jan 2019 = $1666.67 x 0.9434 = $1572.34 or $1572.33

Present Value of $100 Perpetuity at end of year = $100/0.06 = $1666.67
Present Value on Jan 2019 = $1666.67 x 0.89 = $1483.33

Present Value of $100 Perpetuity at end of year = $100x(1.02)/(0.06-0.02) = $2550
Present Value on Jan 2019 = $2550 x 0.9434 = $2405.67 or $2405.66